# Centre for Cyber-Physical Systems IISc, Bengaluru, India

### v1

@@ Abstract
This work aims to synthesize a controller that ensures that an unknown discrete-time system is incrementally input-to-state stable ($\delta$-ISS). In this work, we introduce the notion of $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF), which, in conjunction with the controller, ensures that the closed-loop system is incrementally ISS. To address the unknown dynamics of the system, we parameterize the controller as well as the $\delta$-ISS-CLF as neural networks and learn them by utilizing the sampled data from the state space of the unknown system. To formally verify the obtained $\delta$-ISS-CLF, we develop a validity condition and incorporate the condition into the training framework to ensure a provable correctness guarantee at the end of the training process. Finally, the usefulness of the proposed approach is proved using multiple case studies - the first one is a scalar system with a non-affine non-polynomial structure, the second example is a one-link manipulator system, the third system is a nonlinear Moore-Grietzer model of the jet engine and the final one is a rotating rigid spacecraft model.
@@ 1 Introduction
Unlike traditional stability analysis, which emphasizes convergence to a single equilibrium or nominal trajectory, incremental stability examines the convergence between arbitrary trajectories. Over the years, incremental stability has gained significant attention for its broad applicability, including nonlinear analog circuit modeling [1], cyclic feedback system synchronization [2], symbolic model development [3, 4, 5, 6], stability of interconnected systems [7], oscillator synchronization [8], and complex network analysis [9].
@@
In the course of time, numerous efforts have been made to analyze or verify incremental stability. However, designing controllers that enforce incremental stability has been challenging over the years. Most existing approaches for ensuring incremental stability [18, 19, 20] rely on the assumption of complete knowledge of the dynamics and structure of the system. However, in practice, uncertainties often lead to incomplete or inaccurate system models, rendering such controllers ineffective or poorly performing for the actual system. In general, to address the limitations posed by unknown system dynamics, learning-based methods have gained significant attention. The authors in [21] use the Gaussian process (GP) to collect data from the system and design a backstepping-like controller accordingly under the assumption that the system has a control-affine structure. The newest approach to designing a controller for making a system incrementally ISS [22] also uses a data-driven technique, while the data are collected from the system trajectories. However, this method is only beneficial if the system has a specific polynomial-type structure.
@@
Neural network-based methods have gained prominence for synthesizing controllers alongside Lyapunov or barrier functions. This approach in some existing literature, such as [23, 24, 25, 26] addresses the issue of safety and stability while eliminating the need for explicit knowledge of the system dynamics. Leveraging the universal approximation capabilities of neural networks, functions can be directly synthesized. However, a key challenge is providing formal guarantees, as neural network-based training relies on discrete samples, which cover only a limited part of the continuous state space.
@@
Contributions: Primarily, the work proposes a formally verified neural controller synthesis that ensures incremental input-to-state stability for unknown systems. The previous work of the authors [27] derives the procedure to demonstrate the verification of incremental stability for an unknown system by developing the neural Lyapunov function. In this work, we aim to synthesize the controller such that the closed-loop system exhibits the behavior of incremental stability. To synthesize the controller, we propose the notion of incremental input-to-state stable control Lyapunov function and prove the existence of it under a controller will lead to incremental stability of the closed-loop system. Now, as we use data-driven techniques, we use compactness assumption and the controller is trained using the concept of control barrier function, such that the state space becomes forward invariant. First, we leverage conditions of the Lyapunov function alongside the controller as a robust optimization problem. Since the dynamics of the system is unknown, we utilize data-driven techniques and propose the controller alongside the Lyapunov function to be neural networks and formulate a scenario convex problem corresponding to the robust optimization problem using some Lipschitz continuity assumptions. We construct a dataset from the state space of the system and solve the scenario problem by the verge of training of the neural networks by formulating appropriate loss functions. Following the Lipschitz continuity, we propose a validity condition that gives a formal guarantee that the trained neural controller can make the system incrementally stable irrespective of any initial condition within the state space. Finally, we validate the effectiveness of our approach by applying it to multiple case studies - the first one is a scalar system with a non-affine non-polynomial structure, the second example is a one-link manipulator system, the third system is a nonlinear Moore-Grietzer model of the jet engine and the final one is a rotating rigid spacecraft model.
The key contributions of this paper are highlighted below.
  • 1.
    For the first time, we introduce the notion of incremental input-to-state stable control Lyapunov function.
  • 2.
    Given that the proposed approach requires data for training neural networks, it necessitates reliance on compact sets. To address this, we establish conditions for incremental input-to-state stable control Lyapunov functions under compact sets, ensuring incremental input-to-state stability for system evolving in compact set.
  • 3.
    A novel training framework is presented that simultaneously synthesizes a controller and a verifiably correct incremental input-to-state stable control Lyapunov function, both parameterized as neural networks, for unknown discrete-time systems.
  • 4.
    The previous studies to design the controller needed the system to be fully actuated as well as control affine structure [19, 20]. In this work, we do not require any knowledge of the dynamics and, hence, have synthesized the controller for any general system without the affineness assumption.
  • 5.
    As the controller is a neural network, we can control the system inputs of higher dimensions with a smaller number of external inputs to the controller. The subsection 4.5 details the implication of reduced control input for an incrementally stable system.
The organization of the paper is as follows: In Section 2, we recall the notion of incremental input-to-state stability and introduce the notion of incremental input-to-state stable control Lyapunov function. In Section 3, we propose a learning framework that jointly synthesizes a controller and a verifiably correct incremental input-to-state stable control Lyapunov function, both parametrized as neural networks with the system dynamics being unknown. We pose the problem of synthesizing the controller as an optimization problem and solve it with the help of neural network training. Finally, in section 4, we validate the proposed approach by showcasing multiple case studies and providing a general discussion on the usefulness of the results.
@@ 2 Preliminaries and Problem Formulation
2.1 Notations
The symbols ${\mathbb{N}}$, ${\mathbb{N}}_{0}$, ${\mathbb{R}}$, ${\mathbb{R}}^{+}$, and ${\mathbb{R}}_{0}^{+}$ denote the set of natural, nonnegative integer, real, positive real, and nonnegative real numbers, respectively. A vector space of real matrices with $n$ rows and $m$ columns is denoted by ${\mathbb{R}}^{n\times m}$. A column vector with $n$ rows is represented by ${\mathbb{R}}^{n}$. The Euclidean norm is represented using ||$\cdot$|| ⋅ |. Given a function $\varphi:{\mathbb{N}}_{0}\rightarrow{\mathbb{R}}^{m}$, its sup-norm (possibly $\infty$-norm) is given by $\lVert\varphi\rVert=\sup\{|\varphi(k)|:k\in{\mathbb{N}}_{0}\}$. For $a,b\in{\mathbb{N}}_{0}$ with $a\leq b$, the closed interval in ${\mathbb{N}}_{0}$ is denoted as $[a;b]$. A vector $x\in\mathbb{R}^{n}$ with entries $x_{1},\ldots,x_{n}$ is represented as $[x_{1},\ldots,x_{n}]^{\top}$, where $x_{i}\in\mathbb{R}$ denotes the $i$-th element of the vector and $i\in[1;n]$. A diagonal matrix in ${\mathbb{R}}^{n\times n}$ with non-negative entries is denoted by $\mathcal{D}_{\geq 0}^{n}$. Given a matrix $M\in{\mathbb{R}}^{n\times m}$, $M^{\top}$ represents the transpose of matrix $M$. A continuous function $\alpha:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be class $\mathcal{K}$ if $\alpha(s)>0$ for all $s>0$, strictly increasing and $\alpha(0)=0$. It is class $\mathcal{K}_{\infty}$ if it is class $\mathcal{K}$ and $\alpha(s)\rightarrow\infty$ as $s\rightarrow\infty$. A continuous function $\beta:{\mathbb{R}}_{0}^{+}\times{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a class $\mathcal{KL}$ if $\beta(s,t)$ is a class $\mathcal{K}$ function with respect to $s$ for all $t$ and for fixed $s>0$, $\beta(s,t)\rightarrow 0$ if $t\rightarrow\infty$. For two functions $\gamma,\kappa\in\mathcal{K}_{\infty}$, we say $\gamma<\kappa$ if $\gamma(s)<\kappa(s)$ for all $s>0$. We denote the identity function as $\mathcal{I}_{d}\in\mathcal{K}_{\infty}$. For any compact set $\mathcal{C},\partial\mathcal{C}$ and $int(\mathcal{C})$ denote the boundary and interior of the set $\mathcal{C}$, respectively.
2.2 Incremental Stability
In this work, we consider discrete-time control systems as defined next.
@@ 2.2 Incremental Stability
where $\mathsf{x}(k)\in{\mathbf{X}}$ and $\mathsf{w}(k)\in{\mathbf{W}}$ are the state and external input of the closed-loop system at $k$-th instance, respectively.
@@
The system in (2) is said to be incrementally input-to-state stable ($\delta$-ISS) if there exists a class $\mathcal{KL}$ function $\beta$ and a class $\mathcal{K}_{\infty}$ function $\gamma$, such that for any $k\in{\mathbb{N}}_{0}$, for all $x,\hat{x}\in{\mathbf{X}}$ and any external input sequence $\mathsf{w},\hat{\mathsf{w}}$ the following holds:
$|\mathsf{x}_{x,\mathsf{w}}(k)-\mathsf{x}_{\hat{x},\hat{\mathsf{w}}}(k)|\leq\beta(|x-\hat{x}|,k)+\gamma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert).$ (3)
If $\mathsf{w}=\hat{\mathsf{w}}$, one can recover the notion of incremental global asymptotic stability as defined in [13].
Next, we introduce the notion of incremental input-to-state stable control Lyapunov function ($\delta$-ISS-CLF). The sufficient conditions to ensure incremental stability of the closed-loop system with the controller based on $\delta$-ISS-CLF are defined as follows:
@@
The function $V:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is said to be a $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF) for closed-loop system $\Xi_{g}=({\mathbb{R}}^{n},{\mathbb{R}}^{p},{\mathbb{R}}^{m},f,g)$ in (2), if there exist a controller $g:{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\rightarrow{\mathbb{R}}^{m}$, class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\alpha_{3}$, and a class $\mathcal{K}$ function $\sigma$ such that:
@@
Proof 2.5
The proof follows a similar approach as that of [28, Theorem 1].
@@
Definition 2.6 (Control Forward Invariant Set [29])
A set ${\mathbf{X}}$ is said to control forward invariant with respect to the system (1) if for every initial condition $\mathsf{x}(0)\in{\mathbf{X}}$, there exists a control input $\mathsf{u}(k)\in{\mathbf{U}}$ such that $\mathsf{x}(k)\in{\mathbf{X}},\forall k\in{\mathbb{N}}_{0}$. Now, if there exists a controller $g:{\mathbf{X}}\rightarrow{\mathbf{U}}$ such that the set ${\mathbf{X}}$ becomes forward invariant, then the controller is said to be a forward invariant controller.
Now we introduce the notion of $\delta$-ISS-CLF for the closed-loop system $\Xi_{g}$ where the sets ${\mathbf{X}}\subset{\mathbb{R}}^{n},{\mathbf{W}}\subset{\mathbb{R}}^{p}$ are compact and ${\mathbf{X}}$ is considered to be a control forward invariant under the controller $g$. Then, the definition of $\delta$-ISS-CLF becomes:
@@
The function $V:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is said to be a $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF) for closed-loop system $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$ in (2), where ${\mathbf{X}}$ and ${\mathbf{W}}$ are compact sets, if there exist a forward invariant controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$, class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\alpha_{3}$, and a class $\mathcal{K}$ function $\sigma$ such that:
@@
Proof 2.9
The proof is presented in Appendix A.
@@ 2.3 Control Barrier Function
To ensure that the controller $g$ makes the compact set $X$ forward invariant, we leverage the notion of control barrier function (CBF) introduced next.
@@
Then, $h$ is said to be a control barrier function (CBF) for the system $\Xi$ in (1) if there exists a control input $u\in{\mathbf{U}}$ such that the following condition holds:
$h(f(x,u))-\kappa(h(x))\leq 0,\quad\forall x\in{\mathbf{X}},$ (5)
for some $\kappa\in\mathcal{K}_{\infty}$ with $\kappa\leq\mathcal{I}_{d}$.
Now, based on Definition 2.10, we can design the forward invariant controller $u:=g(x)$ that will make the state-space ${\mathbf{X}}$ control forward invariant. The following lemma allows us to synthesize the controller to enforce control invariance.
@@
The set ${\mathbf{X}}$ will be a control forward invariant set for a system $\Xi$ in (1) if there exists a function $h$ that satisfies the conditions of Definition 2.10.
Proof 2.12
Consider a function $h:{\mathbf{X}}\rightarrow{\mathbb{R}}$ such that it satisfies condition (4), i.e., $h(x)=0,\forall x\in\partial{\mathbf{X}}$ and $h(x)<0,\forall x\in int({\mathbf{X}})$.
Now we assume that there exists a controller $u=g(x)$ such that condition (5) is satisfied. Then, one can easily infer $\forall x\in{\mathbf{X}},h(f(x,g(x)))\leq\kappa(h(x))\leq 0$ which readily implies $f(x,g(x))\in{\mathbf{X}}$. It necessarily implies that starting from any initial condition within the set ${\mathbf{X}}$, the trajectory of the system will remain within the same set ${\mathbf{X}}$ under the action of the controller $g$. Hence, the set ${\mathbf{X}}$ is control forward invariant. This completes the proof.

### v2

@@ Abstract
This work aims to synthesize a controller that ensures that an unknown discrete-time system is incrementally input-to-state stable ($\delta$-ISS). In this work, we introduce the notion of $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF), which, in conjunction with the controller, ensures that the closed-loop system is incrementally ISS. To address the unknown dynamics of the system, we parameterize the controller as well as the $\delta$-ISS-CLF as neural networks and learn them by utilizing the sampled data from the state space of the unknown system. To formally verify the obtained $\delta$-ISS-CLF, we develop a validity condition and incorporate the condition into the training framework to ensure a provable correctness guarantee at the end of the training process. Finally, the usefulness of the proposed approach is proved using multiple case studies.
@@ 1 Introduction
Unlike traditional stability analysis, which emphasizes convergence to an equilibrium or nominal trajectory, incremental stability examines convergence between arbitrary trajectories. Over the years, incremental stability has gained significant attention for its broad applicability, including nonlinear analog circuit modeling [1], cyclic feedback system synchronization [2], symbolic model development [3, 4, 5, 6], oscillator synchronization [7], and complex network analysis [8].
@@
In the course of time, numerous efforts have been made to analyze or verify incremental stability. However, designing controllers that enforce incremental stability has been challenging over the years. Most existing approaches for ensuring incremental stability [17, 18, 19] rely on the assumption of complete knowledge of the dynamics and structure of the system. However, in practice, significant model uncertainties can degrade controller performance, particularly when the assumptions made during design are violated. While classical feedback control offers robustness guarantees to certain perturbations, it may still face limitations under severe uncertainty or high-dimensional nonlinear dynamics. To address such cases, learning-based methods have gained increasing attention. The authors in [20] use the Gaussian process (GP) to collect data from the system and design a backstepping-like controller accordingly under the assumption that the system has a control-affine structure. The newest approach to designing a controller for making a system incrementally ISS [21] also uses a data-driven technique, while the data are collected from the system trajectories. However, this method is only beneficial if the system has a specific polynomial-type structure. Apart from being limited to specific structures, both these works need the system to be identified first before the control design step, which is challenging in all cases.
@@
Neural network-based methods have gained prominence for synthesizing controllers alongside Lyapunov or barrier functions. This approach in some existing literature, such as [22, 23, 24, 25] addresses the issue of safety and stability while eliminating the need for explicit knowledge of the system dynamics. In [26], learning neural Lyapunov functions for stabilizing controllers is discussed, but verification is empirical, whereas [27] designs neural controllers with Lyapunov stability guarantee for state/output feedback under structural assumptions. Leveraging the universal approximation capabilities of neural networks, functions can be directly synthesized. However, a key challenge is providing formal deterministic guarantees, as neural network-based training relies on discrete samples, which cover only a limited part of the continuous state space. While falsification-based approaches [26] have been proposed to mitigate this by adaptively sampling “difficult” regions, they provide only probabilistic confidence; in contrast, we try to enforce deterministic satisfaction of $\delta$-ISS conditions through constrained optimization. The CEGIS-based approach [28] has been a powerful tool as well to verify neural Lyapunov function, but it works in a post-hoc verification scenario, while in this work, we incorporate verification within the training procedure without the need for any other solver.
@@
Contributions: This work proposes a formally verified neural controller synthesis that ensures incremental input-to-state stability for unknown discrete-time systems. Building upon our previous work [29], which verified incremental stability using neural Lyapunov functions, we now focus on controller synthesis that guarantees incremental stability in the closed loop. To this end, we introduce the notion of incremental input-to-state stable control Lyapunov function ($\delta$-ISS-CLF) and prove that its existence under a suitable controller will lead to incremental stability of the closed-loop system. Now, as we use data-driven techniques, we use the compactness assumption, and the controller is trained using the concept of control barrier function, such that the state space becomes forward invariant. First, we formulate the joint synthesis of the controller and Lyapunov function as a robust optimization problem, which is then approximated by a scenario convex program using a finite dataset and considering the controller and Lyapunov function parameterized as neural networks. Following the Lipschitz continuity assumptions, we propose a validity condition that gives a formal guarantee that the trained neural controller can make the system incrementally stable. The proposed method is validated across multiple case studies.
The key contributions of this paper are: $(i)$ For the first time, we introduce the notion of incremental input-to-state stable control Lyapunov function. $(ii)$ Since the proposed approach relies on data for neural network training, we need to work with compact sets. To address this, we establish conditions for $\delta$-ISS-CLF to ensure incremental input-to-state stability for systems evolving in compact sets. $(iii)$ A novel training framework is proposed to jointly synthesize a controller and a verifiably correct $\delta$-ISS-CLF, both realized as neural networks for unknown discrete-time systems. $(iv)$ Unlike previous works that require fully actuated or control-affine systems [18, 19], our approach handles general unknown dynamics.
@@ 2 Preliminaries and Problem Formulation
Notations: The symbols $\mathbb{N}$}, ${\mathbb{N}}_{0}$, $\mathbb{R}$}, $\mathbb{R}$$^{+}$, and ${\mathbb{R}}_{0}^{+}$ denote the set of natural, nonnegative integers, real, positive real, and nonnegative real numbers, respectively. The vector space of real matrices with $n$ rows and $m$ columns is denoted by ${\mathbb{R}}^{n\times m}$. The set of column vectors with $n$ rows is represented by ${\mathbb{R}}^{n}$. The Euclidean norm is represented using $|\cdot|$. Given a function $\varphi:{\mathbb{N}}_{0}\rightarrow{\mathbb{R}}^{m}$, its sup norm is given by $\lVert\varphi\rVert=\sup\{|\varphi(k)|:k\in{\mathbb{N}}_{0}\}$. For $a,b\in{\mathbb{N}}_{0}$ with $a\leq b$, the closed interval in ${\mathbb{N}}_{0}$ is denoted by $[a;b]$. A vector $x\in\mathbb{R}^{n}$ with entries $x_{1},\ldots,x_{n}$ is represented as $x_{1},\ldots,x_{n}]^{\top}$, where $x_{i}\in\mathbb{R}$ denotes the $i$-th element of the vector and $i\in[1;n]$. The set of diagonal matrices with non-negative entries in ${\mathbb{R}}^{n\times n}$ is denoted by $\mathcal{D}_{\geq 0}^{n}$. Given a matrix $\in{\mathbb{R}}^{n\times m}, MM^{\top}$ represents the transpose of matrix $M$. A continuous function $\alpha:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be class $\mathcal{K}$ if $\alpha(s)>0$ for all $s>0$, strictly increasing and $\alpha(0)=0$. It is class $\mathcal{K}_{\infty}$ if it is class $\mathcal{K}$ and $\alpha(s)\rightarrow\infty$ as $s\rightarrow\infty$. A continuous function $\beta:{\mathbb{R}}_{0}^{+}\times{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a class $\mathcal{KL}$ if $\beta(s,t)$ is a class $\mathcal{K}$ function with respect to $s$ for all $t$ and for fixed $s>0$, $\beta(s,t)\rightarrow 0$ if $t\rightarrow\infty$. For two functions $\gamma,\kappa\in\mathcal{K}_{\infty}$, we say $\gamma<\kappa$ if $\gamma(s)<\kappa(s)$ for all $s>0$. For any compact set $\mathcal{C},\partial\mathcal{C}$ and $int(\mathcal{C})$ denote the boundary and interior of the set $\mathcal{C}$, respectively.
2.1 Incremental Input-to-State Stability
Consider the discrete-time control systems as defined next.
@@ 2.1 Incremental Input-to-State Stability
where $\mathsf{x}(k)\in{\mathbf{X}}$ and $\mathsf{w}(k)\in{\mathbf{W}}$ are the state and external input of the closed-loop system at $k$-th instance, respectively. Note that both $f$ and $g$ are considered to be Lipschitz continuous.
@@
The system $\Xi_{g}=({\mathbb{R}}^{n},{\mathbb{R}}^{p},{\mathbb{R}}^{m},f,g)$ in (2) is said to be incrementally input-to-state stable ($\delta$-ISS) if there exist a class $\mathcal{KL}$ function $\beta$ and a class $\mathcal{K}_{\infty}$ function $\gamma$, such that for any $k\in{\mathbb{N}}_{0}$, for all $x,\hat{x}\in{\mathbf{X}}\subseteq{\mathbb{R}}^{n}$ and any external input sequence $\mathsf{w},\hat{\mathsf{w}}$ (where $\mathsf{w},\hat{\mathsf{w}}:{\mathbb{N}}\rightarrow{\mathbf{W}}$) the following holds:
[…]
If $\mathsf{w}=\hat{\mathsf{w}}=0$ and ${\mathbf{X}}={\mathbb{R}}^{n}$, one can recover the notion of incremental global asymptotic stability as defined in [12].
Next, we define the incremental input-to-state stable control Lyapunov function ($\delta$-ISS-CLF) and present the sufficient conditions ensuring incremental stability of the closed-loop system under a $\delta$-ISS-CLF–based controller.
@@
The function $V:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{{\mathbb{R}}_{0}^{+}}$ is said to be a $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF) for closed-loop system $\Xi_{g}=({\mathbb{R}}^{n},{\mathbb{R}}^{p},{\mathbb{R}}^{m},f,g)$ in (2), if there exist a controller $g:{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\rightarrow{\mathbb{R}}^{m}$, class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\alpha_{3}$, and a class $\mathcal{K}$ function $\sigma$ such that:
@@
Proof:
The proof is similar to that of [30, Theorem 1]. $\square$
@@
Definition 2.5 (Robustly Forward Invariant Set [31])
A set $\mathbf{X}$} is said to be robustly forward invariant with respect to the system (2) if for every $(x,w)\in{\mathbf{X}}\times{\mathbf{W}}$, there exists some control input $u:=g(x,w)\in{\mathbf{U}}$ such that $f(x,g(x,w))\in{\mathbf{X}}$. Now, if there exists a controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ such that the set $\mathbf{X}$} becomes robustly forward invariant, then the controller is said to be a forward invariant controller corresponding to the robustly invariant set $\mathbf{X}$} with respect to external input $\mathbf{W}$}.
Now we introduce the notion of $\delta$-ISS-CLF for the closed-loop system $\Xi_{g}$ where the sets ${\mathbf{X}}\subset{\mathbb{R}}^{n},{\mathbf{W}}\subset{\mathbb{R}}^{p}$ are compact and $\mathbf{X}$} is considered to be a robustly forward invariant under the controller $g$. Then, the definition of $\delta$-ISS-CLF becomes:
@@
The function $V:{\mathbf{X}}\times{\mathbf{X}}\rightarrow{\mathbb{R}}_{0}^{+}$} is said to be a $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF) for closed-loop system $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$ in (2), where $\mathbf{X}$} and $\mathbf{W}$} are compact sets, if there exist a forward invariant controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$, class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\alpha_{3}$, and a class $\mathcal{K}$ function $\sigma$ s.t.
@@
Proof:
The proof is presented in Appendix A. $\square$
@@ 2.2 Control Barrier Function
The notion of CBF is introduced to ensure that the compact set $\mathbf{X}$} becomes robustly forward invariant. Since the $\delta$-ISS-CLF in Definition 2.6 is defined over a compact set, it is essential to ensure that the closed-loop trajectories remain within $\mathbf{X}$} for the validity of the incremental stability guarantees. We leverage the notion of control barrier function (CBF) introduced next.
@@
Then, $h$ is said to be a control barrier function (CBF) for the system $\Xi$ in (2) if for every $(x,w)\in{\mathbf{X}}\times{\mathbf{W}}$, there exists some control input $u:=g(x,w)\in{\mathbf{U}}$ such that the following holds:
${h(f(x,g(x,w)))-h(x)\leq 0.}$ (5)
Now, based on Definition 2.8, we can design the forward invariant controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ that will make the state-space $\mathbf{X}$} control forward invariant. The following lemma allows us to synthesize the controller to enforce control invariance.
@@
Consider a dt-CS $\Xi$ in (1) and let $h:{\mathbf{X}}\rightarrow{\mathbb{R}}$ be a control barrier function as defined in Definition 2.8. Then, the controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ satisfying condition (5) will make the set $\mathbf{X}$} robustly forward invariant.
Proof:
Consider a function $h:{\mathbf{X}}\rightarrow{\mathbb{R}}$ such that it satisfies condition (4), i.e., $h(x)=0,\forall x\in\partial{\mathbf{X}}$ and $h(x)<0,\forall x\in int({\mathbf{X}})$. Now we assume that there exists a controller $g(x,w)$ such that condition (5) is satisfied. Then, one can easily infer $\forall x\in{\mathbf{X}},h(f(x,g(x,w)))\leq h(x)\leq 0$ which readily implies $f(x,g(x,w))\in{\mathbf{X}}$ which implies that the trajectory of the system will remain within the same set $\mathbf{X}$} under the action of the controller $g$. Hence, the set $\mathbf{X}$} is control forward invariant. $\square$